# KTWIN: A Serverless Kubernetes-based Digital Twin Platform

### v1

@@ 6 Conclusion
To further enhance the capabilities of our platform, future research and development efforts will address some of the limitations identified and new requirements that were not in the initial proposed design. The lack of scalability in dispatcher components, representing a bottleneck for the system, will be addressed by implementing multiple consumers for the same event type published to the Event Broker. The routing limitations of MQTT messages using Cloud Events can be further analyzed using MQTT 5 headers and improvements in the routing rules of the Event Broker. In addition, future research includes the evaluation of KTWIN in Edge deployment scenarios and how to distribute the control and application plane components in case the solution is deployed for multiple landscapes. Finally, future work can include requirements of the Digital Twin Capabilities Table (CPT) not tackled in this study such as enabling event store data integration for existing and external enterprise systems and applications or implementing services use cases for advanced analytics and Machine Learning.
@@ Acknowledgments
This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001.

### v3

@@ 7 Conclusion
To further enhance the capabilities of our platform, future research and development efforts will address some of the limitations identified and new requirements that were not in the initial proposed design. The lack of scalability in dispatcher components, representing a bottleneck for the system, will be addressed by implementing multiple consumers for the same event type published to the Event Broker. In addition, future research includes the evaluation of KTWIN in Edge deployment scenarios and how to distribute the control and application plane components in case the solution is deployed to Cloud and Edge locations. Future works can also evaluate the modeling and deployment of different Digital Twin use cases scenarios such as Manufactoring, Energy, Healthcare or other Smart Cities scenarios. Finally, future works can include Digital Twin Capabilities Table (CPT) requirements not tackled in this study such as provide advanced 3D visualization of domain-related data in charts, graphs or dashboards, enable Data Analysis and Analytics using collected event data, implement use cases levering Machine Learning and Edge AI foundation requirements within the platform, and research and validate requirements related to data governance and data privacy. The implemented prototype and deployment setup are available at https://github.com/Open-Digital-Twin.
@@ Acknowledgments
This research is part of the INCT of Intelligent Communications Networks and the Internet of Things (ICoNIoT) funded by CNPq Process 405940/2022-0 and CAPES Finance Code 88887.954253/2024-00. This study was also partially funded by CAPES Finance Code 001.